If a borrower borrows $3,000 at 6%, single payment in one year. If the loan has a $100 origination fee on the loan, what is the total amount they would have to pay? What is the annual percent rate then?
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Amount borrowed                  3,000.00
Origination fee                   (100.00)
Amount received on borrowing                  2,900.00
Interest on Loan = 3000 *6%                     180.00
Total Amount to be paid = 180+ 3000                  3,180.00
Annual percent rate = (3180 - 2900)/3180 8.81%